# New flexible versions of extended generalized Pareto distribution for count data

### v3

@@ 1 Introduction
Many statistical models are available for analyzing non-negative count data. Standard frameworks include the Poisson distribution, which suits count data with equal mean and variance, and the Negative Binomial distribution, which addresses overdispersion by allowing variance to exceed the mean. For data with an excessive number of zeros, ZI models such as the ZI Poisson (ZIP) and ZI Negative Binomial (ZINB) extend these distributions to account for additional zero inflation. The DGPD, on the other hand, is tailored for modeling discrete exceedances above a high threshold, making it ideal for analyzing extreme values and rare events. Each model addresses different data characteristics and analysis needs, providing versatile tools for statistical analysis. In practice, however, count data are often overdispersed to these distributions, and sometimes extreme observations are also present. For instance, the count data of upheld complaints of automobile insurance companies complaints rankings in New York City and the number of doctors visits to the hospital are heavy-tailed datasets and have excessive zeros. In both datasets, 47.53% and 42% of the values are zeros, indicating no upheld complaints registered and no recorded doctor visits at the hospital, respectively. Conversely, 6.18% and 2.48% of the observations are extreme values in both datasets, respectively, identified using the right fence with $c=3$ of the Tukey fence method (Schwertman et al.,, 2004). We fitted ZINB to both datasets and constructed their quantile-quantile (Q-Q) plots. Figure 1 (a, b) shows that the model fits well to the lower and central parts of the data distribution but exhibits deficiencies at the upper tail. Consequently, building a model for the count data sets with ZI and extreme observations becomes a significant problem and has not yet been properly studied.
@@
Modeling discrete exceedances is a viable approach to analyzing extreme tail behavior, offering valuable insights into rare events. The peak-over-threshold (POT) method, as introduced by Pickands, (1975), commonly approximates the distribution of exceedances above a high threshold to the Generalized Pareto Distribution (GPD). Several adaptations of the GPD for discrete data are available in the literature (Krishna and Pundir,, 2009, Buddana and Kozubowski,, 2014, Kozubowski et al.,, 2015). More recently, Hitz et al., (2024) discussed the DGPD for modeling the tail behavior of integer-valued random variables. Consequently, when the threshold is sufficiently high, standard extreme value models, such as the POT method with DGPD approximation (Ranjbar et al.,, 2022, Hitz et al.,, 2024), can be applied to model the exceedances. Detailed information is provided in Section 2. Choosing an appropriate threshold $u$ is critical for effectively applying the POT approximation. Setting the threshold too low can introduce bias into the estimates, as the DGPD is justified only in an asymptotic sense. On the other hand, setting the threshold too high reduces the number of data points, increasing the estimation variance. In practice, selecting a suitable threshold in a continuous setting often involves parameter stability plots and mean residual life plots, as suggested by Coles, (2001, ch. 4) and Davison and Smith, (1990). A suitable threshold can also be identified where the mean of the exceedances (mean residual life) starts to exhibit a linear relationship with the threshold value (Daouia et al.,, 2023). In the discrete POT approach, defining a threshold at a high quantile remains necessary, which poses challenges due to the discrete nature of the data.
@@
Selecting an appropriate threshold is often challenging in practice, as stability plots and mean residual life plots may not always clearly indicate the best threshold. Furthermore, thresholds that are straightforward for practitioners to interpret might be too low to accurately estimate extremes. When applying the POT model, this can lead to modeling discrete exceedances above a threshold that is too low from the standpoint of extreme value theory, thereby invalidating the approximation of the DGPD. An alternative method for modeling discrete exceedances above a low threshold involves using a discrete version of one of the three extended generalized Pareto models introduced by Papastathopoulos and Tawn, (2013).
@@
In this contribution to the literature, we introduce new versions of the DGPD to enhance the model proposed by Ahmad et al., (2024). The proposed models exhibit flexibility in three key areas: First, effectively model the entire range of non-negative integer-valued data, accommodating a wide variety of count data scenarios; second, correctly handle ZI integer-valued data, providing accurate and reliable results even when datasets contain a significant proportion of zero counts; and, third, model the exceedances above a low threshold without requiring a censoring in the likelihood, simplifying the modeling process in cases where setting an appropriate high threshold is challenging. These enhancements offer a comprehensive and flexible approach to modeling discrete data, particularly in situations where existing models fail, see Figure 1.
@@ 2 Modeling framework
where $(a)_{+}=\max(a,0)$, $\beta>0$, and $-\infty<\xi<+\infty$ represent the scale and shape parameters of the distribution, respectively.
As the threshold increases, the limiting distribution of observations converges to a GPD as explained in Theorem 1
Theorem 1.
For a random variable $X$ that takes values in the interval in $[0,x_{\mathcal{F}})$ where $x_{\mathcal{F}}\in(0,\infty)\cup\{\infty\}$, suppose that there exists a strictly positive sequence $a_{u}$ such that the distribution of $a_{u}^{-1}(X-u)|X\geq u$ weakly converge to a non-degenerate probability distribution on $[0,\infty)$ as $u\rightarrow x_{\mathcal{F}}$, then this distribution is the GPD (Balkema and De Haan,, 1974). Thus, for sufficiently large $u$,
@@
However, applying the GPD to discrete data can be problematic, as noted by Hitz et al., (2024). They suggested that for an integer-valued random variable $Y$, the tail distribution can be better approximated by discretizing the CDF (1) for large $u$.
@@
with $\beta>0$ and $\xi\geq 0$. The distribution given in (2) is known as the DGPD and has been explored in the literature from different perspectives; see, for instance, Hitz et al., (2024, and references therein).
A notable drawback of the DGPD is its reliance on observations exceeding a certain threshold, which imposes an artificial dichotomy on the data (i.e., values are classified as either above or below the threshold). This threshold selection can be complex, particularly with datasets containing many tied values. In contrast, some continuous extreme value models aim to avoid the need of threshold selection by modeling the entire range of data (Frigessi et al.,, 2002, Carreau and Bengio,, 2009, MacDonald et al.,, 2011, Papastathopoulos and Tawn,, 2013, Naveau et al.,, 2016, Stein,, 2021)
A noteworthy extension to GPD proposed by Papastathopoulos and Tawn, (2013) incorporates an additional shape parameter without affecting the tail behavior. Additional shape parameter stabilized the threshold estimation for the GPD parameter, allowing a lower threshold to be chosen. In addition to modeling the lower tail and the bulk of the distribution, Naveau et al., (2016) identified two conditions to ensure compliance with the EVT. The transition between the two tails on $[0,1]$ can take various forms. We follow the footsteps of Naveau et al., (2016) to develop a model for an entire range of discrete data. They use the idea of integral transformation to simulate GPD random draws, i.e. $\mathcal{F}_{\beta,\xi}^{-1}(U),$ where $U\sim\mathcal{U}(0,1)$ represents a uniformly distributed random variable on $(0,1)$ and $\mathcal{F}_{\beta,\xi}^{-1}$ denotes the inverse of the CDF (1). Based on this, we can construct a family of distributions for the random variables
$Z=\mathcal{F}_{\beta,\xi}^{-1}\left(\mathcal{G}^{-1}(U)\right),$ (3)
where $\mathcal{G}$ is a CDF on $[0,1]$ and $U\sim\mathcal{U}(0,1)$. Clearly the CDF of $Z$ is $\mathcal{G}(\mathcal{F}(z;\beta,\xi))$. The key problem is to find a function $\mathcal{G}$, which preserves the upper tail behavior with shape parameter $\xi$ and also controls the lower tail behavior. Naveau et al., (2016) defined restrictions for validity of $\mathcal{G}$ functions. For instance, the tail of $\mathcal{G}$ denoted by $\bar{\mathcal{G}}=1-\mathcal{G}$ has to satisfy
@@
[…]
Definition 3.
Let $Y$ be an integer-valued random variable. The probability mass function (pmf) of DEGPD is defined by using definition 1 as
$\Pr(Y=y)=\mathcal{G}\left(\mathcal{F}\left({y+1};\beta,\xi\right);\kappa\right)-\mathcal{G}\left(\mathcal{F}\left({y};\beta,\xi\right);\kappa\right),\qquad y\in\mathbb{N}_{0},$ (6)
where the parameter $\kappa$ belongs to $\mathcal{G}$ family, $\beta$ and $\xi$ are usual parameters of GPD. The explicit formula of CDF of DEGPD is
$\Pr(Y\leq y)=\mathcal{G}\left(\mathcal{F}\left({y+1};\beta,\xi\right),\kappa\right),$ (7)
and quatntile function is
$q_{p}=\left\{\begin{array}[]{ll}\left\lceil{\frac{\beta}{\xi}\left\{\left(1-\mathcal{G}^{-1}(p)\right)^{-\xi}-1\right\}}\right\rceil-1,&\mbox{if }\xi>0\\ \\ \left\lceil{-\beta\log\left(1-\mathcal{G}^{-1}(p)\right)}\right\rceil-1,&\mbox{if }\xi=0\\ \end{array}\right.$ (8)
We define another pmf by following Lambert, (1992) for the case of zero-inflation in count data.
Definition 4.
Suppose $Y$ is observed with excessive zeros relative to those observed under the (6). In that case, the pmf with additional $0\leq\pi\leq 1$ parameter, which represent zero values, is defined as
$\Pr(Y=y)=\left\{\begin{array}[]{ll}\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(1;\beta,\xi\right);\kappa\right)&y=0\\ (1-\pi)\left[\mathcal{G}\left(\mathcal{F}\left(y+1;\beta,\xi\right);\kappa\right)-\mathcal{G}\left(\mathcal{F}\left(y;\beta,\xi\right);\kappa\right)\right]&y=1,2,\ldots\end{array}\right.$ (9)
$\Pr(Y\leq y)=\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(y+1,\beta,\xi\right),\kappa\right),\qquad y\in\mathbb{N}_{0},$ (10)
and the quantile function is derived
$q_{p^{*}}=\left\{\begin{array}[]{lc}\lceil\frac{\beta}{\xi}\left[\left\{1-\mathcal{G}^{-1}(p^{*})\right\}^{-\xi}-1\right]\rceil-1,&\xi>0\\ \lceil{-\beta\log\left\{1-\mathcal{G}^{-1}(p^{*})\right\}}\rceil-1,&\xi=0,\\ \end{array}\right.$ (11)
with $10<p<$ 1 and $10<p^{*}=(p-\pi)/(1-\pi)<$ 1.
In this paper, we use three parametric expressions for $\mathcal{G}(\cdot)$ that fulfill the conditions (4) and conditions $(C_{1},C_{2}$ and $C_{3})$ presented in Gamet and Jalbert, (2022). The following three models fulfill the above-mentioned conditions and are deemed suitable for modeling the bulk of the distribution while maintaining the stability of heavier tail behavior.
@@
$\mathcal{G}(u;\kappa)=u^{\kappa}$, $\kappa>0$;
@@
$\mathcal{G}(u;\kappa)=\frac{2}{2\Phi\left(\sqrt{\kappa}\right)-1}\Big{[}\Phi\left(\sqrt{\kappa}(u-1)\right)-\left\{1-\Phi(\sqrt{\kappa})\right\}\Big{]}$, the defined $\mathcal{G}(u)$ is the cdf of truncated normal distribution with precision $\kappa>0$ and $\Phi$ is the cdf of standard normal distribution. After satisfying the conditions $(C_{1},C_{2}$ and $C_{3})$, the density of function of truncated normal distribution for variable $U$ is defined as
$f_{u}(u;\kappa)=\left\{\begin{array}[]{ll}\frac{2\sqrt{\kappa}}{2\Phi\left(\sqrt{\kappa}\right)-1}\phi\left(\sqrt{\kappa}(u-1)\right)&\text{if}\hskip 14.22636pt\kappa>0\\ 1&\text{if}\hskip 14.22636pt\kappa\rightarrow 0,\end{array}\right.$
@@
We consider the truncated Beta distribution as another choice to add more flexibility in the density of $U$. Let $D_{x}(\alpha_{1},\alpha_{2})$ indicates incomplete beta function:
[…]
@@
$Beta(u;\alpha_{1},\alpha_{2})=\frac{\Gamma(\alpha_{1}+\alpha_{2})}{\Gamma(\alpha_{1})\Gamma(\alpha_{2})}u^{\alpha_{1}-1}(1-u)^{\alpha_{2}-1},$
with $00\leq u\leq 1,\alpha_{1}>$ 0 and $\alpha_{2}>0$.The distribution with parameters $(\kappa,\kappa)$ bounded on interval $(\omega,1/2)$ with $\kappa>0$ and $20<\omega<1/$ 2
$f_{u}(u;\kappa)=\frac{1}{D_{1/2}(\kappa,\kappa)-D_{\omega}(\kappa,\kappa)}Beta\left((1/2-\omega)u+\omega;\kappa,\kappa\right).$ (12)
The expression (12) rescaled to unit interval and satisfied the conditions $(C_{1},C_{2}$ and $C_{3})$. For more details, we refer to see Gamet and Jalbert, (2022). The cdf of $U$ corresponding to pmf defined in (12) is written as
[…]

### v4

@@ 1 Introduction
Many statistical models are available for analyzing non-negative count data. Standard frameworks include the Poisson distribution, which suits count data with equal mean and variance, and the Negative Binomial distribution, which addresses overdispersion by allowing variance to exceed the mean. For data with an excessive number of zeros, ZI models such as the ZI Poisson (ZIP) and ZI Negative Binomial (ZINB) extend these distributions to account for additional zero inflation. The DGPD, on the other hand, is tailored for modeling discrete exceedances above a high threshold, making it ideal for analyzing extreme values and rare events. Each model addresses different data characteristics and analysis needs, providing versatile tools for statistical analysis. In practice, however, count data are often overdispersed to these distributions, and sometimes extreme observations are also present. For instance, the count data of upheld complaints of automobile insurance companies complaints rankings in New York City and the number of doctors visits to the hospital are heavy-tailed datasets and have excessive zeros. In both datasets, 47.53% and 42% of the values are zeros, indicating no upheld complaints registered and no recorded doctor visits at the hospital, respectively. Conversely, 6.18% and 2.48% of the observations are extreme values in both datasets, respectively, identified using the right fence with $c=3$ of the Tukey fence method [21]. We fitted ZINB to both datasets and constructed their quantile-quantile (Q-Q) plots. Figure 1 (a, b) shows that the model fits well to the lower and central parts of the data distribution but exhibits deficiencies at the upper tail. Consequently, building a model for the count data sets with ZI and extreme observations becomes a significant problem and has not yet been properly studied.
@@
Modeling discrete exceedances is a viable approach to analyzing extreme tail behavior, offering valuable insights into rare events. The peak-over-threshold (POT) method, as introduced by Pickands [19], commonly approximates the distribution of exceedances above a high threshold to Generalized Pareto Distribution (GPD). Several adaptations of the GPD for discrete data are available in the literature [14, 4, 13]. More recently, Hitz et al. [12] discussed the DGPD for modeling the tail behavior of integer-valued random variables. Consequently, when the threshold is sufficiently high, standard extreme value models, such as the POT method with DGPD approximation [20, 12], can be applied to model the exceedances. Detailed information is provided in Section 2. Choosing an appropriate threshold $u$ is critical for effectively applying the POT approximation. Setting the threshold too low can introduce bias into the estimates, as the DGPD is justified only in an asymptotic sense. On the other hand, setting the threshold too high reduces the number of data points, increasing the estimation variance. In practice, selecting a suitable threshold in a continuous setting often involves parameters stability plots and mean residual life plots, as suggested by Coles [7, ch. 4] and Davison and Smith [9]. A suitable threshold can also be identified where the mean of the exceedances (mean residual life) starts to exhibit a linear relationship with the threshold value [8]. In the discrete POT approach, defining a threshold at a high quantile remains necessary, which poses challenges due to the discrete nature of the data.
@@
Selecting an appropriate threshold is often challenging in practice, as stability plots and mean residual life plots may not always clearly indicate the best threshold. Furthermore, thresholds that are straightforward for practitioners to interpret might be too low to accurately estimate extremes. When applying the POT model, this can lead to modeling discrete exceedances above a threshold that is too low from the standpoint of extreme value theory, thereby invalidating the approximation of the DGPD. An alternative method for modeling discrete exceedances above a low threshold involves using a discrete version of one of the three extended generalized Pareto models introduced by Papastathopoulos and Tawn [18].
@@
In this contribution to the literature, we introduce new versions of the DGPD to enhance the model proposed by Ahmad et al. [1]. The proposed models exhibit flexibility in three key areas: First, effectively model the entire range of non-negative integer-valued data, accommodating a wide variety of count data scenarios; second, correctly handle ZI integer-valued data, providing accurate and reliable results even when datasets contain a significant proportion of zero counts; and, third, model the exceedances above a low threshold without requiring a censoring in the likelihood, simplifying the modeling process in cases where setting an appropriate high threshold is challenging. These enhancements offer a comprehensive and flexible approach to modeling discrete data, particularly in situations where existing models fail, see Figure 1.
@@ 2 Modeling framework
where $(a)_{+}=\max(a,0)$, $\beta>0$, and $\infty<\xi<+\infty$ represent the scale and shape parameters of the distribution, respectively.
As the threshold increases, the limiting distribution of observations converges to a GPD as explained below.
For a random variable $X$ that takes values in the interval in $0,x_{\mathcal{F}})$ where $x_{\mathcal{F}}\in(0,\infty)\cup\{\infty\}$, suppose that there exists a strictly positive sequence $a_{u}$ such that the distribution of $a_{u}^{-1}(X-u)|X\geq u$ weakly converges to a non-degenerate probability distribution on $0,\infty)$ as $u\rightarrow x_{\mathcal{F}}$, then this distribution is the GPD [3]. Thus, for sufficiently large $u$,
@@
However, applying the GPD to discrete data can be problematic, as noted by Hitz et al. [12]. They suggested that for an integer-valued random variable $Y$, the tail distribution can be better approximated by discretizing the CDF (1) for large $u$.
@@
$\Pr(Y-u=k|Y\geq u)=\mathcal{F}(k+1;\beta,\xi)-\mathcal{F}(k;\beta,\xi),\hskip 28.45274ptk\in\mathbb{N}_{0},$ (2)
with $\beta>0$ and $\xi\geq 0$, the $\mathcal{F}(.;\beta,\xi)$ is the CDF of GPD defined in (1). The distribution given in (2) is known as the DGPD and has been explored in the literature from different perspectives; see, for instance, Hitz et al. [12, and references therein].
A notable drawback of the DGPD is its reliance on observations exceeding a certain threshold, which imposes an artificial dichotomy on the data (i.e., values are classified as either above or below the threshold). This threshold selection can be complex, particularly with datasets containing many tied values. In contrast, some continuous extreme value models aim to avoid the need of threshold selection by modeling the entire range of data [10, 5, 16, 18, 17, 23]
A noteworthy extension to GPD proposed by Papastathopoulos and Tawn [18] incorporates an additional shape parameter without affecting the tail behavior. Additional shape parameter stabilized the threshold estimation for the GPD parameter, allowing a lower threshold to be chosen. In addition to modeling the lower tail and the bulk of the distribution, Naveau et al. [17] identified two conditions to ensure compliance with the EVT. The transition between the two tails on $[0,1]$ can take various forms. We follow the footsteps of Naveau et al. [17] to develop a model for an entire range of discrete data. They use the idea of integral transformation to simulate GPD random draws, i.e. $\mathcal{F}_{\beta,\xi}^{-1}(U),$ where $U\sim\mathcal{U}(0,1)$ represents a uniformly distributed random variable on $(0,1)$ and $\mathcal{F}_{\beta,\xi}^{-1}$ denotes the inverse of the CDF (1). Based on this, we can construct a family of distributions for the random variables
$Z=\mathcal{F}_{\beta,\xi}^{-1}\left(\mathcal{G}^{-1}(U)\right),$ (3)
where $\mathcal{G}$ is a CDF on $[0,1]$ and $U\sim\mathcal{U}(0,1)$. Clearly the CDF of $Z$ is $\mathcal{G}(\mathcal{F}(z;\beta,\xi))$. The key problem is to find a function $\mathcal{G}$, which preserves the upper tail behavior with shape parameter $\xi$ and also controls the lower tail behavior. Naveau et al. [17] defined restrictions for validity of $\mathcal{G}$ functions. For instance, the tail of $\mathcal{G}$ denoted by $\bar{\mathcal{G}}=1-\mathcal{G}$ has to satisfy
$\lim_{\nu\to 0}\frac{\bar{\mathcal{G}}(1-\nu)}{\nu}$ $=$ $a,\mbox{ for some finite $a>0$ (upper tail behavior),}$ (4)
$\lim_{\nu\to 0}\frac{\mathcal{G}(\nu)}{\nu^{\kappa}}$ $=$ $c,\mbox{ for some finite $c>0$ (lower tail behavior).}$ (5)
Definition 3.
Let $Y$ be an integer-valued random variable. The probability mass function (PDF) of DEGPD is defined by using definition 1 as
$\Pr(Y=k)=\mathcal{G}\left(\mathcal{F}\left({k+1};\beta,\xi\right);\kappa\right)-\mathcal{G}\left(\mathcal{F}\left({k};\beta,\xi\right);\kappa\right),\qquad k=0,1,2,\ldots,$ (6)
where the parameter $\kappa$ represent the lower tail of the distribution which modeled by $\mathcal{G}$ defined in Models (i)-(iii) later, $\beta$ and $\xi$ are usual parameters of GPD. The $\mathcal{G}\left(\mathcal{F}\left({.};\beta,\xi\right);\kappa\right)$ is the CDF of EGPD in accordance with Models (i)-(iii). For $\xi>0$, we have shown that the EGPD CDF satisfies extreme value theory properties, exhibiting a regularly varying tail. Moreover, the model introduced in (6) is also regularly varying, sharing the same tail behavior as in the continuous case, and belongs to the maximum domain of attraction. Detailed proofs of the EGPD and DEGPD properties for each Model (i)–(iii) are provided in the Supplementary Material. The explicit formula of the CDF of DEGPD is
$\Pr(Y\leq k)=\mathcal{G}\left(\mathcal{F}\left({k+1};\beta,\xi\right),\kappa\right),$ (7)
and quatntile function is
@@
We define another PDF by following Lambert [15] for the case of zero-inflation in count data.
Definition 4.
Suppose $Y$ is observed with excessive zeros relative to those observed under the (6). In that case, the PDF with additional $0\leq\pi\leq 1$ parameter, which represent zero values, is defined as
$\Pr(Y=y)=\left\{\begin{array}[]{ll}\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(1;\beta,\xi\right);\kappa\right)&k=0\\ (1-\pi)\left[\mathcal{G}\left(\mathcal{F}\left(k+1;\beta,\xi\right);\kappa\right)-\mathcal{G}\left(\mathcal{F}\left(k;\beta,\xi\right);\kappa\right)\right]&k=1,2,\ldots\end{array}\right$. (9)
$\Pr(Y\leq k)=\pi+(1-\pi)\mathcal{G}\left(\mathcal{F}\left(k+1,\beta,\xi\right),\kappa\right),\qquad k=0,1,2,\ldots,$ (10)
and the quantile function is derived
@@
with $0<p<1$ and $0<p^{*}=(p-\pi)/(1-\pi)<1$.
In this paper, we use three parametric expressions for $\mathcal{G}(\cdot)$ that fulfill the conditions (4) and conditions $(C_{1},C_{2}$ and $C_{3})$ presented in Gamet and Jalbert [11]. The following three models fulfill the above-mentioned conditions and are deemed suitable for modeling the bulk of the distribution while maintaining the stability of heavier tail behavior.
@@
$\mathcal{G}(\nu;\kappa)=\nu^{\kappa}$, $\kappa>0$;
@@
$\mathcal{G}(\nu;\kappa)=\frac{2}{2\Phi\left(\sqrt{\kappa}\right)-1}\Big[\Phi\left(\sqrt{\kappa}(\nu-1)\right)-\left\{1-\Phi(\sqrt{\kappa})\right\}\Big]$, the defined $\mathcal{G}(\nu)$ is the CDF of truncated normal distribution with precision $\kappa>0$ and $\Phi$ is the CDF of standard normal distribution. After satisfying the conditions $(C_{1},C_{2}$ and $C_{3})$, the density of function of truncated normal distribution for variable $U$ is defined as
$f_{\nu}(\nu;\kappa)=\left\{\begin{array}[]{ll}\frac{2\sqrt{\kappa}}{2\Phi\left(\sqrt{\kappa}\right)-1}\phi\left(\sqrt{\kappa}(\nu-1)\right)&\text{if}\hskip 14.22636pt\kappa>0\\ 1&\text{if}\hskip 14.22636pt\kappa\rightarrow 0,\end{array}\right$.
@@
We consider the truncated Beta distribution as an alternative choice to add more flexibility to the density of $U$. Let $D_{x}(\alpha_{1},\alpha_{2})$ indicates incomplete beta function:
$D_{x}(\alpha_{1},\alpha_{2})=\int_{0}^{x}t^{\alpha_{1}-1}(1-t)^{\alpha_{2}-1}dt,$
@@
$\text{Beta}(\nu;\alpha_{1},\alpha_{2})=\frac{\Gamma(\alpha_{1}+\alpha_{2})}{\Gamma(\alpha_{1})\Gamma(\alpha_{2})}\nu^{\alpha_{1}-1}(1-\nu)^{\alpha_{2}-1},$
with $0\leq\nu\leq 1,\alpha_{1}>0$ and $\alpha_{2}>0$.The distribution with parameters $(\kappa,\kappa)$ bounded on interval $(\omega,1/2)$ with $\kappa>0$ and $0<\omega<1/2$
$f_{\nu}(\nu;\kappa)=\frac{1}{D_{1/2}(\kappa,\kappa)-D_{\omega}(\kappa,\kappa)}Beta\left((1/2-\omega)\nu+\omega;\kappa,\kappa\right).$ (12)
The expression (12) rescaled to the unit interval and satisfied the conditions $(C_{1},C_{2}$ and $C_{3})$. For more details, we refer to see, for instance, Gamet and Jalbert [11]. The CDF corresponding to the PDF defined in (12) is written as
$\mathcal{G}(\nu;\kappa)=\left\{\begin{array}[]{ll}0&\text{for}\hskip 14.22636pt\nu<0;\\ \frac{D_{\left\{(1/2-\omega)\nu-\omega\right\}}(\kappa,\kappa)-D_{\omega}(\kappa,\kappa)}{D_{1/2}(\kappa,\kappa)-D_{\omega}(\kappa,\kappa)}&\text{for}\hskip 14.22636pt0\leq\nu\leq 1;\\ 1&\text{for}\hskip 14.22636pt\nu>1.\end{array}\right$. (13)